OpenStudy (anonymous):

the product of a number and 6 less than the number is 72. Find the number

OpenStudy (anonymous):

12 times 6 is 72

OpenStudy (anonymous):

what does product mean and can you show me an example.

OpenStudy (anonymous):

when you multiply two numbers the answer will be the product of the two

OpenStudy (anonymous):

12 * 6 = 72

OpenStudy (anonymous):

ok there is two answers -12 or 6 and -6 or 12 which one is it

OpenStudy (across):

Mathematically, call this number\[a.\]Then\[a(a-6)=72,\]\[a^2-6a-72=0,\]\[a=12\]or\[a=-6.\]Both of these work.

OpenStudy (anonymous):

both because u have the same numbers just the sign is different... but a negative times a positive is always a negative number